OpenStudy (anonymous):

Mario answered 75% of the 40 problems on his history homework correctly. How many problems on his homework did he answer correctly? A 19 B 30 C 28 D 10

OpenStudy (anonymous):

\(.75\times 40\)

OpenStudy (anonymous):

or if you prefer, \(\frac{3}{4}\times 40\)

OpenStudy (anonymous):

.75 x 40

OpenStudy (anonymous):

30 ?